TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-02-26 22:15:00
subject: 2\08 Pt-1 ISS On Orbit Status 08-02-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
2\08 ISS On-Orbit Status 08-02-2003
Part 1 of 2

10 Feb 2003

ISS On-Orbit Status 2/8/03

"We cannot stress enough how blessed and honored we feel to be counted 
as members of the NASA family. We proudly support the noble goals and
objectives of NASA and we will continue to support NASA in its finest 
and its darkest hours. It is our deepest hope that you also will 
continue to share in our belief and support of NASA's dreams." (Public 
statement [excerpt] by the husbands, wives, and children of the NASA 
Astronaut Corps, 2/7/03.)

All ISS systems continue to function nominally, except as noted 
previously or below. Day 77 in space for the Expedition 6 crew, after 
a very tough week for all, but especially so for the Columbia crew's 
families and also for Ken, Don and Nikolai in their high outpost.

The station crew had their regular off-duty Saturday, except for 
weekend routine tasks, necessary repair work and daily maintenance.

After wake-up at the regular 1:00am EST, morning inspection, morning 
hygiene and breakfast, the ISS residents completed the weekly 3-hr. 
house cleaning.  As part of today's housekeeping, Bowersox performed 
head cleaning on VTR-2 (video tape recorder #2), which is getting 
close to the regular 250-hr.  servicing. [The "uborka stantsii" 
focuses on removal of food waste products, cleaning of compartments 
with vacuum cleaner, wet cleaning of surfaces with disinfectants and 
cleaning of fan screens to avoid temperature rises.]

FE-1 Nikolai Budarin completed the periodic inspection of the BRPK
air/condensate separator of the SRVK water processing system. [The 
BRPK separator contains porous cermet hydrophilic ("water attracting") 
tubes through which the gas-liquid mixture from the heat exchanger 
moves. They separate the air from the condensate, but when the 
separator exceeds its service life or malfunctions, incomplete 
separation of the atmospheric condensate occurs, and the water then 
collects under the "sheet" of porous fluoroplastic. This is the main 
focus of the regular inspection.]

Budarin, supported by FE-2/SO Don Pettit, then removed and replaced a 
pipe conduit at the BPK condensate micropump (MR2-2G) of the BRPK 
separator unit.  [The pipe conduit connects the BPK MR2-2G to a 
solenoid valve on the BPK air line. Objective of the 2-hr. 
installation of the new conduit, which is fitted with a check valve 
and control valves, is to protect the micropump from moisture 
ingestion and decrease the flow of supply air, thereby mitigating the 
adverse effect that the operation of the BPK unit has had on the life 
span of the BRPK air/condensate separators.]

On MCC-M Go-ahead, Budarin also terminated the regeneration cycle on
adsorbent bed #1 of the BMP harmful impurities filter unit, returning 
it to Purify mode. Later, after #1 was determined to work properly, he 
switched channel #2 to the 24-hr bake-out process for regeneration.

At 4:00am and 5:50am, FE-1 Budarin conducted two downlink sessions of 
the video records of the TEKh-20 Plasma Crystal experiment (PK-3) 
conducted on 1/22. [After playback of the videocassette, Nikolai left 
all Russian Telescience components connected for the next downlinks, 
currently scheduled for 2/12 (next Wednesday).]

Nikolai also performed the task-list item of reviewing his upcoming 
next session with the MBI-8 biomedical "Profilaktika" (preventive 
health maintenance) fitness session, scheduled for 2/10-12. [During 
these three days, Budarin will use the VELO ergometer on Day 1, the 
Load Trainer on Day 2 and the TVIS treadmill on Day 3. The test is 
similar to the MedOps MO-3 test, except that it uses the TEEM-100M gas 
analyzer during exercise operations, blood lactate determination and 
subjective load assessment.]

Don Pettit again collected the two regular daily carbon dioxide (CO2)
readings in the SM and Lab with the U.S. CDMK (CO2 monitoring kit).

The daily routine inspection/maintenance of the SOSh life support 
system, incl. the ASU toilet system, was completed by CDR Ken 
Bowersox, while Pettit prepared the IMS inventory system for database 
auto import/export. [After database update files are uplinked, the 
crew needs to auto import them into the system].

Budarin conducted his daily 24-hr. checkup and watering of the Russian 
BIO-5 Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment.

For "Diatomeya" world ocean observations, on the Russian task list, 
Budarin today was to observe and image seawater bloom features (TsKO) 
and cloud structures in the waters of the Atlantic, Pacific, and 
Indian Oceans.  [Bioproductive areas were identified in the Atlantic 
west of South Africa, in the Benguela Current and Upwelling and along 
the south boundary of the Gulf Stream, in the Pacific in the frontal 
portion of the Peruvian Current.]

Today's targets for the Russian "Uragan" (hurricane) earth observation
experiment, also a task list item using the Nikon D1 digital still 
camera (f80-400 zoom lens), were the Himalayas (convergence shooting, 
i.e., a series of consecutive frames while keeping the target at the 
center of the camera's field of view), the city of Burgas and the 
Panama Canal.

At 7:05am EST, attitude control was handed over to the Russian MCS 
(motion control system) thrusters of the Service Module for the 
dynamic test of the US-21 matching unit installed in the Progress 10P 
yesterday. [The test firings of Progress thruster manifolds 1 and 2 
were to be completed at 10:48am, and control authority handed back to 
U.S. CMGs (control moment thrusters) at 11:00am to resume momentum 
management in LVLH (local vertical/local horizontal). Progress 
thruster yaw and pitch control was incorporated into the SM's 
propulsion control algorithm.]

PFCs (private family conferences) were scheduled for Ken Bowersox and 
Don Pettit today and are timelined for all crewmembers tomorrow.

Yesterday's Velocicalc airflow measurements determined low air flow 
from the Node aft port and Lab aft port fans. [To support 
troubleshooting, the crew was asked tocheckthe position of the two 
Node air damper valves. Also, the ground will deactivate and 
reactivate each suspect fan to determine the run and 
startupcharacteristics. As a comparison, the starboard aft fan will be 
likewisebe deactivated and activated.Further troubleshooting will 
follow next week but as of now there is no other action required on 
the crew's part.]

MSG (microgravity science glovebox) continues inoperative. 
Troubleshooting (t/s) efforts are underway at POC (Payload Operations 
Center) by ESA. [In order to minimize the risk of additional problems 
associated with closing a circuit breaker into a possible shorted 
circuit, the t/s will wait until ESA and MSFC have completed a 
detailed audit of the as-built MSG configuration, analyzed the failure 
data, and re-reviewed ground test data. Initial t/s procedures are 
expected to be available in the next few days, and some time has been 
scheduled on the crew's timeline early next week to allow direct 
discussion with the MSG ground team. A detailed information package on 
MSG activities has also been uplinked.]

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.